alinson has a jelly beans that sell for $0.75 per pound and small eggs that sells for $2.25 per pound. how many pouinds of each should she use to obtain 12 pounds of an easter mix that sells for $1.75 per pound?
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

2.25x+.75(12-x)=1.75*12
2.25x+9-.75x=21
1.5x=21-9
1.5x=12
x=12/1.5
x=8 pounds of small eggs are used.
12=8=4 pounds of jelly beans were used.
Proof:
2.25*8+.75*4=1.75*12
18+3=21
21=21

Answer by stanbon(75887)

alinson has a jelly beans that sell for $0.75 per pound and small eggs that sells for $2.25 per pound. how many pouinds of each should she use to obtain 12 pounds of an easter mix that sells for $1.75 per pound?
------------------------------------------------------------
Equation:
Value + Value = Value
75x + 225(12-x) = 175*12
75x + 225*12 - 225x = 175*12
-150x = -75*12
x = 6 lbs (amt of 75cent jelly beans in the mix)
12-x = 6 lbs (amt of 225cent small eggs in the mix)
=====================================================
Cheers,
Stan H.